OpenStudy (anonymous):

3. What is the length of the segment joining the points at (–3, 0) and (5, 6)? _________ units 4. What is the length of the segment joining the points at (1, –8) and (5, 0)? Round to the nearest tenth if necessary. _________ units

OpenStudy (anonymous):

@jdoe0001 @TheSmartOne

OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ -3}}\quad ,&{\color{blue}{ 0}})\quad % (c,d) &({\color{red}{ 5}}\quad ,&{\color{blue}{ 6}})\\ % (a,b) &({\color{red}{ 1}}\quad ,&{\color{blue}{ -8}})\quad % (c,d) &({\color{red}{ 5}}\quad ,&{\color{blue}{ 0}}) \end{array}\qquad % distance value d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)

OpenStudy (jdoe0001):

so their "distance" is also their "length

OpenStudy (jdoe0001):

well. the length of the segment between them for that matter

OpenStudy (anonymous):

I'm still confused, can you put it any simpler?

OpenStudy (anonymous):

@jdoe0001 @TheSmartOne@jim_thompson5910 @e.mccormick @Data_LG2 @dan815

TheSmartOne (thesmartone):

You plug in the points into the formula and solve for the distance.

TheSmartOne (thesmartone):

So for right now, lets consider these 2 points \(\sf (3, 0) ~and ~(5, 6)\) So using the formula \(\Large\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\) And the points are in the form of \(\sf(x_1,y_1),(x_2,y_2)\)

TheSmartOne (thesmartone):

And so then \(\sf x_1 = 3 \\ x_2 =5 \\y_1=0 \\y_2=6\) And then just plug it in to the huge distance formula :)

OpenStudy (anonymous):

Can you help me fill in the formula?

TheSmartOne (thesmartone):

\(\large\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\) \(\sf x_1 = 3 \\ x_2 =5 \\y_1=0 \\y_2=6\) \(\large\sqrt{(5-3)^2+(6-0)^2}\)

OpenStudy (anonymous):

2√10?

TheSmartOne (thesmartone):

Correct :)

TheSmartOne (thesmartone):

For the second question :P \(\large\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\) \(\sf x_1 = 1 \\ x_2 =5 \\y_1=-8 \\y_2=0\) \(\large\sqrt{(5-1)^2+(0-(-8))^2}\)

OpenStudy (anonymous):

But for the first one how many units would that be?

TheSmartOne (thesmartone):

It would just be \( 2\sqrt{10}\) units

OpenStudy (anonymous):

But I can put that symbol in...

OpenStudy (anonymous):

1. 2√10 2. 4√5

TheSmartOne (thesmartone):

That is correct. If it doesn't allow you to put that symbol.. then we could simplify that and then 1)6.325 2)8.944 That is what you would get if you plugged it into the calculator and rounded upto 3 digits.

OpenStudy (anonymous):

Thank you. :)

TheSmartOne (thesmartone):

Np :)